OpenStudy (anonymous):

Ramon is playing a game in which he must pull two blocks out of a bag containing red and yellow blocks. He cannot look, and he cannot replace the block. The bag has 4 more red blocks than yellow blocks. a) Write and simplify an expression that represents Ramon's probability of picking a red block, then a yellow block. b) What is the probability that Ramon pulls a red block then a yellow block if there are 6 yellow blocks in the bag before his pick? c) What is the probability that Ramon pulls two yellow blocks if there are 6 yellow blocks in the bag before his first pick?

OpenStudy (anonymous):

if the number of yellow blocks is n, then the number of red blocks is n+4 and the total number of blocks is n+n+4=2n+4 probability you pick a red block is \[\frac{n+4}{2n+4}\] and then the probability you pick a yellow block given the first is red is \[\frac{n}{2n+3}\] multiply these together and get \[\frac{(n+4)n}{(2n+4)(2n+3)}\]

OpenStudy (anonymous):

if there are six yellow blocks in the bag then there are ten red ones and the total is 16

OpenStudy (anonymous):

probability if red then yellow is \[\frac{10}{16}\times \frac{6}{15}\]